OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. (2 points) Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28 Select one: a. Equation 1 has more solutions than equation 2. b. Equation 1 and Equation 2 have the same number of solutions. c. Equation 2 has more solutions than Equation 1. d. The number of solutions cannot be determined.

OpenStudy (anonymous):

HALP!!!!!!!!!!!!!!!

OpenStudy (anonymous):

\[41+28=69\]

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

c

OpenStudy (anonymous):

ok

OpenStudy (radar):

Solve equation 1, count the number of solutions. Solve equation 2, count the number of solutions. I will solve equation 2, step by step for you, you then solve 1 and compare the number of solutions.

OpenStudy (radar):

Equation 2: |2x +13| =28 To solve you must let the value of 28 take on both signs. Solving for +28, 2x + 13 = 28, 2x = 28 -13, 2x = 15 giving us x=15/2 Solving for -28, 2x +13 = -28, 2x = -28 - 13, 2x =-41 or x = -41/2 Equation 2 has two solutions.

OpenStudy (radar):

Oh by the way choice c is incorrect.

OpenStudy (radar):

@slobo_da_hobo Oh, by the way choice c. is incorrect.

OpenStudy (anonymous):

@radar, so what would be the answer?

OpenStudy (radar):

Solve equation 1 the same way as I showed you step by step, it is simple algebra. Let the value 41 be both a + and a - 41. solve for each value and you will get the number of solutions so that you can compare with the number of solutions to equation. Remember,there were two solutions for equation 2.

OpenStudy (radar):

I will start you off. Solve for 5x + 6 = 41 Solve for 5x + 6 = -41 Just solve for x in each of those and count the number of values you get for x

OpenStudy (radar):

As you know I can give you the answer, that is something you are learning how to do.